# Multitask Learning with Stochastic Interpolants

### v1

@@ 1.1 Related works
Our framework offers a unified approach to inverse problems, contrasting with traditional methods that require problem-specific variational optimization procedures Pereyra et al. (2015). Recent diffusion-based approaches Chung et al. (2023); Song et al. (2022); Kawar et al. (2022) typically need guided sampling trajectories tailored to each task. Similarly, methods using MCMC/SMC sampling Coeurdoux et al. (2024); Sun et al. (2024), variational approximations Mardani et al. (2023); Alkan et al. (2023), or optimization-driven techniques Wang et al. (2024); Daras et al. (2022) remain fundamentally task-specific. Our approach encodes solution paths within the interpolant operator structure itself, enabling multiple inverse problems to be addressed through appropriate operator path selection during inference—without additional training. Our work can also be seen as a way to formalize methods that seek to clean data corrupyed in various ways Bansal et al. (2022)
@@ 3.3 Fine-tuning and posterior sampling
involving data $x_{1}^{r}$ from the posterior distribution. {pinkbox} {restatable}propositiontwoeta Let
@@
This proposition is proven in Appendix A as a corollary of Proposition A that relates the probability distribution of $I_{r}$ to that of $I$. Proposition 15 offers a way to sample the posterior distribution without retraining, by using the drifts (18) and (19) in the ODE (7) or the SDE (8).
@@ 5.1 Multitask inpainting and sequential generation
CelebA and AFHQ-Cat. We present benchmark results for all methods across various image restoration tasks, evaluating the average peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) on 100 test images from each dataset: AFHQ-Cat ($256\times 256$) and CelebA ($128\times 128$). To assess the performance of our methodology, we employed two types of masking: square masks of sizes $40\times 40$ and $80\times 80$ with added Gaussian noise of standard deviation 0.05, and random masks covering $70\%$ of image pixels with Gaussian noise of standard deviation 0.01.
@@
Table 1: PSNR and SSIM metrics for image inpainting methods on AFHQ-Cat and CelebA datasets.
Method AFHQ-Cat CelebA
@@
OT-ODE 28.36 0.865 28.84 0.914 28.84 0.838 23.88 0.874
D-Flow 33.07 0.938 29.70 0.893 31.37 0.888 26.69 0.833
Flow-Priors 32.33 0.945 29.40 0.858 31.76 0.909 25.85 0.822
PnP-Flow 33.54 0.953 30.59 0.943 32.98 0.930 26.87 0.904
@@ 5.2 Posterior sampling in the $\phi^{4}$-model
The additional term plays the role of a reward. The results of the generation based on Proposition 15 are shown on Figure 5, which indicate that our approach permits to valid sample configurations (as verified by their magnetization) of the posterior without retraining.
@@ Appendix A Proofs
[Posterior distributions]propositiontwomu Let $\mu_{\alpha,\beta}$ and $\mu^{r}_{\alpha,\beta}$ be the probability distributions of the stochastic interpolants defined in (14) and (15), respectively. Assume that $\alpha$ and $\beta$ are invertible, that the equations (20) for $\alpha_{r}$, $\beta_{r}$ in Proposition 15 have a solution, and that $x_{r}$ is given by (21). Then these distributions are related, up to a constant independent of $x$ and $x_{r}$, as
@@
If in this equation we replace $\alpha$ by $\alpha_{r}$, $\beta$ by $\beta_{r}$, $x$ by $x_{r}$, and $\phi(x)$ by $\phi(x)e^{R(\alpha,\beta,x)}$, and multiply both side by $|\alpha_{r}|/|\alpha|$ it becomes
@@ Appendix C Additional experimental results
Figure 8: Additional images demonstrating the inpainting task: Random inpainting is shown in the top panels, while block inpainting is displayed in the bottom panels. The left panels depict images from the AFHQ dataset, and the right panels show images from the CelebA dataset

### v2

@@ 1.1 Related works
Our framework offers a unified approach to inverse problems, contrasting with traditional methods that require problem-specific variational optimization procedures Pereyra et al. (2015). Recent diffusion-based approaches Chung et al. (2023); Song et al. (2022); Kawar et al. (2022); Martin et al. (2025) typically need guided sampling trajectories tailored to each task. Similarly, methods using MCMC/SMC sampling Coeurdoux et al. (2024); Sun et al. (2024), variational approximations Mardani et al. (2023); Alkan et al. (2023), or optimization-driven techniques Wang et al. (2024); Daras et al. (2022) remain fundamentally task-specific. Our approach encodes solution paths within the interpolant operator structure itself, enabling multiple inverse problems to be addressed through appropriate operator path selection during inference—without additional training. Our work can also be seen as a way to formalize methods that seek to clean data corrupyed in various ways Bansal et al. (2022)
@@ 3.3 Fine-tuning and posterior sampling
involving data $x_{1}^{r}$ from the posterior distribution.
{pinkbox}{restatable}
propositiontwoeta Let
@@
This proposition is proven in Appendix A as a corollary of Proposition A that relates the probability distribution of $I_{r}$ to that of $I$. Proposition 3.3 offers a way to sample the posterior distribution without retraining, by using the drifts (18) and (19) in the ODE (7) or the SDE (8).
@@ 5.1 Multitask inpainting and sequential generation
CelebA and AFHQ-Cat. We present benchmark results for all methods across various image restoration tasks, evaluating the average peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) on 100 test images from each dataset: AFHQ-Cat ($256\times 256$) and CelebA ($128\times 128$). To assess the performance of our methodology, we employed two types of masking: square masks of sizes $40\times 40$ and $80\times 80$ with added Gaussian noise of standard deviation 0.05, and random masks covering $70\%$ of image pixels with Gaussian noise of standard deviation 0.01. We benchmark our method against four state-of-the-art interpolant-based restoration methods: PnP-Flow Martin et al. (2025), Flow-Priors Zhang et al. (2024), D-Flow Ben-Hamu et al. (2024), OT-ODE Pokle et al. (2024).
@@
Table 1: PSNR and SSIM metrics for image inpainting methods on CelebA and AFHQ-Cat datasets.
Method CelebA AFHQ-Cat
@@
Pokle et al. (2024) 28.36 0.865 28.84 0.914 28.84 0.838 23.88 0.874
Ben-Hamu et al. (2024) 33.07 0.938 29.70 0.893 31.37 0.888 26.69 0.833
Zhang et al. (2024) 32.33 0.945 29.40 0.858 31.76 0.909 25.85 0.822
Martin et al. (2025) 33.54 0.953 30.59 0.943 32.98 0.930 26.87 0.904
@@ 5.2 Posterior sampling in the $\phi^{4}$-model
The additional term plays the role of a reward. The results of the generation based on Proposition 3.3 are shown on Figure 5, which indicate that our approach permits to valid sample configurations (as verified by their magnetization) of the posterior without retraining.
@@ Appendix A Proofs
[Posterior distributions]propositiontwomu Let $\mu_{\alpha,\beta}$ and $\mu^{r}_{\alpha,\beta}$ be the probability distributions of the stochastic interpolants defined in (14) and (15), respectively. Assume that $\alpha$ and $\beta$ are invertible, that the equations (20) for $\alpha_{r}$, $\beta_{r}$ in Proposition 3.3 have a solution, and that $x_{r}$ is given by (21). Then these distributions are related, up to a constant independent of $x$ and $x_{r}$, as
@@
If in this equation we replace $\alpha$ by $\alpha_{r}$, $\beta$ by $\beta_{r}$, $x$ by $x_{r}$, and $\phi(x)$ by $\phi(x)e^{R(\alpha,\beta,x)}$, and multiply both side by $|\alpha_{r}|/|\alpha|$ it becomes:
@@ Appendix C Additional experimental results
Figure 8: Additional images demonstrating the inpainting task: Random inpainting is shown in the top panels, while block inpainting is displayed in the bottom panels. The left panels depict images from the AFHQ dataset, and the right panels show images from the CelebA dataset.